# On Bureau’s classification of quadratic differential equations in two variables free of movable critical points

### v2

@@ 1. Introduction
Within the study of ordinary differential equations in the complex domain, the following problem is attributed to Goursat: to classify the systems of equations
@@
with $P$ and $Q$ rational in $y$ and $z$, and analytic in the independent variable $t$, which are free of movable critical points, with two systems differing by a transformation which is analytic in the independent variable and birational in the dependent ones considered equal [Gar60]. Following Painlevé, a critical point of a function is a singular point around which at least two of its branches are permuted (see the first of his Leçons de Stockholm [Pai97]). The above is a natural and broad generalization of the problem solved by Painlevé and Gambier concerning second-order equations in one variable, which lead to the discovery of the Painlevé transcendents. Garnier wondered if such general systems could define transcendents beyond those that appear in Painlevé and Gambier’s classification, and estimated this to be a difficult problem. One may restrict it by considering systems (1) with $P$ and $Q$ polynomial in $y$ and $z$. These polynomial equations already form a rich class, which naturally contains Painlevé’s first two, and which, as shown by Malmquist [Mal23, §33], contains the other Painlevé equations as well (see also [Oka79]). Goffar-Lombet investigated the question for polynomial systems having some particular homogeneous polynomials, of degrees four and five, as highest-degree terms, obtaining precise and complete results for such systems [GL74]. In 1980, Kimura and Matuda announced a classification of polynomial systems of the form (1) free of movable critical points, of degree up to five [KM80], which does not seem to have had a follow-up from their part. Shortly afterwards, in a series of articles published in the Bulletin de la Classe des Sciences of the Académie Royale de Belgique ([Bur81a], [Bur81b], [Bur81c]), Bureau classified quadratic systems up to affine transformations. His results are summarized in Table III in [Bur81c, §28] (he uses the term stable to refer to systems that are free of movable critical points). He later studied higher degree polynomial systems in the subsequent articles of the series. The aim of this article is to revisit and revise Bureau’s classification of the quadratic case:
@@
with $\xi$ affine in $\zeta$ and holomorphic in $t$, and $\tau$ a biholomorphism onto its image, such that the system solved by $\xi$, as a function of $\tau$, is one of those appearing in Table 1.
Reciprocally, all systems in Table 1 are free of movable critical points. They are integrated either in an elementary way or via some Poincaré transcendent (also indicated in Table 1).
@@
   with a condition of order $n-1$ (Section 3.6)
[…]
@@
Table 1. Representatives of the quadratic systems free of movable critical points (Theorem 1). Here, $n\in\mathbf{Z}$, $a$, $b$, etc. are functions, $H$ and $K$ are constants.
System Equations Int.
@@
   with $f^{\prime\prime}=0$, $(p^{\prime\prime}-p^{3}-fp)^{\prime}=0$
P. II
@@
P. I
Table 1 (continued).
[…]
Some systems belonging to different rows of Table 1 are birationally equivalent, or are birationally equivalent to some Painlevé equation. The following result will be proved in the course of establishing Theorem 1.
@@
We have the following birational equivalences.
  1. (1)
    A system
    • IX.B(2) with parameter $q$; or
    • IX.B(5) with parameter $q$; or
    • XIV with parameters $(r,p)$, for $q=\frac{1}{12}(p^{\prime}+p^{2}-r)$;
@@
    is equivalent to the system V with parameter $f=q^{\prime\prime}-6q^{2}$ by a bimeromorphic map $\phi:\mathbf{C}^{2}\times\mathbf{C}\dashrightarrow\mathbf{C}^{2}\times\mathbf{C}$ of the form $(\zeta,t)\mapsto(\xi(\zeta,t),t)$, with $\xi$ birational as a function of $\zeta$ for most values of $t$. (System V is the first-order system associated to the Painlevé I equation and some of its degenerations.)
  2. (2)
    A system
    • XIII, with parameters $(f,p)$, for $\alpha=2p^{3}+fp-p^{\prime\prime}$; or
@@
    • IX.B(3), with parameters $(f,H)$, for $\alpha=H-\frac{1}{2}f^{\prime}$;
@@
    is birationally equivalent, by a bimeromorphic map as above, to the cubic first-order system associated to the equation
    $y^{\prime\prime}=2y^{3}+fy+\alpha\quad(f^{\prime\prime}=0,\;\alpha\in\mathbf{C}),$
    giving the Painlevé II equation and some of its degenerations.
  3. (3)
    A system XII, with parameters $(f,H,K)$, for $\alpha=\frac{1}{2}H-K-f^{\prime}$ and $\beta=-H^{2}$, is birationally equivalent, by a bimeromorphic map as above, to the rational first-order system associated to the equation
    $y^{\prime\prime}=\frac{(y^{\prime})^{2}}{2y}+\frac{3}{2}y^{3}+4fy^{2}+2(f^{2}-\alpha)y-\frac{\beta^{2}}{2y}\quad(f^{\prime\prime}=0,\;\alpha,\beta\in\mathbf{C}),$ (2)
    giving the Painlevé IV equation and some of its degenerations.
@@
The less satisfactory part of the classification concerns system VIII, the only one for which the necessary and sufficient conditions for the absence of movable critical points do not explicitly appear, neither in Bureau’s work, nor in Table 1. These systems form an infinite family, indexed by a positive integer $n$, and the complexity of a certain necessary condition for the absence of movable critical points grows with $n$; it seems difficult to give an explicit expression for this condition for every $n$. At the same time, the equations are, conceptually, very simple, as they factor through Riccati equations. In any case, there remains the question of whether the above necessary condition is also a sufficient one. We will show that this is indeed the case, by studying the birational geometry of these systems, in the spirit of Okamoto’s work on the spaces of initial conditions of the Painlevé equations  [Oka79], which is, on its turn, motivated by Painlevé’s 17th of his Leçons de Stockholm [Pai97]. This will be done in Section 3.6.
@@
There is no system X in Table 1. Table III in [Bur81c] contains a “system X,” which summarizes the results in [Bur81b, §15]. There, under a different normalization, the analysis in [Bur81b, §14], the one leading to system VI, is duplicated. Bureau uses this to establish the connections with other equations that he had previously studied, but this does not contribute towards a classification.
@@ 2.1. Quadratic homogeneous equations and univalence
The limit as $\alpha\to 0$ is the autonomous, homogeneous equation $\mathrm{d}\xi/\mathrm{d}\tau=P_{n}(\xi,t_{0})$. If the original equation is free of movable critical points, the latter will have only single-valued solutions [Pai00] (see also [Bur80]). Thus, the first problem, in our setting, is to determine the equations of the form
@@ V $\begin{array}[]{l}y^{\prime}=0\\ z^{\prime}=6y^{2}\end{array}$
VII $\begin{array}[]{l}y^{\prime}=0\\ z^{\prime}=yz\end{array}$infty
@@
$\begin{array}[]{l}y^{\prime}=y[y-(n+2)z]\\ z^{\prime}=-z(ny+z)\end{array}$ $11,n+1,-n-$ 1
@@ XI $\begin{array}[]{l}y^{\prime}=y(y-z)\\ z^{\prime}=z(z-y)\end{array}$ $2,2,\infty$
XII $\begin{array}[]{l}y^{\prime}=y(2z+y)\\ z^{\prime}=-z(2y+z)\end{array}$ $3,3,3$
@@ XIV $\begin{array}[]{l}y^{\prime}=y(2z-y)\\ z^{\prime}=z(y-z)\end{array}$ $2,3,6$
Table 2. Representatives of the univalent quadratic homogeneous systems (Bureau’s normalization), with the indices of their simple radial orbits
@@
$-\frac{s^{\prime\prime}}{(s^{\prime})^{2}}=F(s),$ (4)
for some rational function $F$ (a similar equation is satisfied by $1/s$). Briot and Bouquet gave necessary and sufficient conditions on $F$ for the univalence of a function satisfying such an equation [BB55]: at the finite values of $s$, $F$ has only simple poles, and their residues are either equal to $-1$, or are of the form $11/k-$ 1 for $k\in\mathbf{Z}$, plus the analogous condition at $s=\infty$ coming from the equation satisfied by $1/s$ (see also Painlevé [Pai00, §10]; see [Gui06, Section 2.2] for an interpretation in terms of affine structures). The index of the pole of (4) is the associated integer $k$ (we will say that a simple pole with residue $-1$ has index $\infty$). We refer the reader to the closing remarks in [Con99, §2.3] for a discussion around this terminology.
A radial orbit for an equation of the form (3) is a line through $(0,0)$ that is invariant by the equation (that is a union of solutions). If not every line through the origin is a radial orbit, the equation has three of them (counted with multiplicity). The poles of the associated equation (4) come from radial orbits, although some radial orbits may fail to produce them. The index of a radial orbit is the one of the corresponding pole in equation (4); a simple radial orbit which produces no poles for (4) is said to have index $1$. If an equation has three radial orbits, with indices $n_{i}$,
@@
The classification of the equations (3) with three different radial orbits amounts to solving this equation as a Diophantine one (admitting that $n_{i}$ can take the value $\infty$), from which the last six lines of Table 2 are obtained (the linear equivalence class of such an equation is determined by its indices, and the equations thus obtained happen to be univalent). The other lines of this table follow from organizing quadratic homogeneous vector fields according to the multiplicities of their radial orbits, to the indices of the simple ones, and from imposing the conditions guaranteeing the simplicity of the poles of the associated equation (4).

### v3

@@ Abstract.
Published as: A. Guillot, On Bureau’s Classification of Quadratic Differential Equations in Two Variables Free of Movable Critical points, Stud. Appl. Math. 155 no. 2 (2025), 70083, doi:10.1111/sapm.70083
[…]
@@ 1. Introduction
[…]
For a complex-valued function of one complex variable, a direct source of multivaluedness is the existence of critical points, singular points around which at least two of its branches are permuted (see [Pai02]). For the solutions of differential equations, these points may be separated into fixed (occurring at fixed values of the independent variable for all solutions of a given equation) and movable (appearing at values of the independent variable depending upon the particular solution of the equation). In order to tackle the previous problem, Painlevé proposed the study of a wider class of differential equations, those whose solutions are free of movable critical points (“équations à points critiques fixes”). This property acquired a relevance of its own, and is sometimes referred to as the Painlevé Property. He saw this class as a natural extension of that of linear differential equations, and developed tools for its study. His methods evidenced the utility of considering equations whose dependence upon the independent variable is only analytic (and not necessarily algebraic).
Painlevé and Gambier’s efforts lead to a classification of the second-order equations of the form $y^{\prime\prime}=R(y^{\prime},y,t)$, with $R$ rational in $y$ and $y^{\prime}$ and analytic in the independent variable $t$, which are free of movable critical points: up to fractional linear transformations in $y$, depending analytically on $t$, and reparametrizations of the independent variable, they belong to some fifty explicit classes of equations that are either integrated in an elementary way, or are one of six model equations, the Painlevé equations, labeled I trough VI, which cannot be reduced to either linear or first-order equations [Pai02], [Gam10] (see also [Inc44, Ch. XIV], [Bur64], [Hil97, Ch. 12.2]). Painlevé anticipated the applicability of these functions, and his insight has been proven amply correct, as they “have arisen in a variety of important physical applications including statistical mechanics, random matrices, plasma physics, nonlinear waves, quantum gravity, quantum field theory, general relativity, nonlinear optics and fibre optics” [Cla06] (see also [AC91, §7.1.6], [GLS02, Ch. 10]).
The results of Painlevé and Gambier have been generalized to higher-order and higher-degree equations in one variable; we refer to [CM20, App. A] for a presentation of some of these generalizations.
One may also consider the problem of classifying systems of differential equations of the form
@@
with $P$ and $Q$ rational in $y$ and $z$, and analytic in the independent variable $t$, which are free of movable critical points. The problem solved by Painlevé and Gambier is a particular case of it. In the above problem, it is natural to consider equivalence up to changes of coordinates which are analytic in the independent variable and birational in the dependent ones, as these changes of coordinates preserve both the form of the equations and the property of being free of movable critical points. Garnier, who attributes the problem to Goursat, wondered if such general systems could define transcendents (functions) beyond those that appear in Painlevé and Gambier’s classification, and estimated this to be a difficult problem [Gar60]. One may restrict it by considering systems of the form (1) with $P$ and $Q$ polynomial in $y$ and $z$, up to either affine or birational equivalence (the first preserves the class of systems of a given degree; the second, the systems whose solutions are free of movable critical points). These polynomial equations already form a rich class, which naturally contains Painlevé’s first two, and which, as shown by Malmquist [Mal23, §33], contains the other Painlevé equations as well (see also [Oka79]). Goffar-Lombet investigated the question for polynomial systems having some particular homogeneous polynomials, of degrees four and five, as highest-degree terms, obtaining precise and complete results for them [GL74]. In 1980, Kimura and Matuda announced a classification of polynomial systems of the form (1) free of movable critical points, of degree up to five [KM80], which does not seem to have had a follow-up from their part. Shortly afterwards, in a series of articles published in the Bulletin de la Classe des Sciences of the Académie Royale de Belgique ([Bur81a], [Bur81b], [Bur81c]), Bureau classified quadratic systems up to affine transformations. His results are summarized in Table III in [Bur81c, §28] (he uses the term stable to refer to systems that are free of movable critical points). The aim of this article is to revisit and revise Bureau’s classification:
@@
with $\xi$ affine in $\zeta$ and holomorphic in $t$, and $\tau$ a biholomorphism onto its image, such that the system solved by $\xi$, as a function of $\tau$, is the restriction to $\mathbf{C}^{2}\times\phi(V)$ of one of those appearing in Table 1.
Reciprocally, all systems in Table 1 are free of movable critical points. They are integrated either in an elementary way or via some Painlevé transcendent (also indicated in Table 1).
@@
   with a condition of order n1n-1 (Section 3.6)
$\text{IX.A}_{0}(n)$
@@
   with $f^{\prime\prime}=0$, $(p^{\prime\prime}-2p^{3}-fp)^{\prime}=0$
Painlevé II
@@
Painlevé I
Table 1. Representatives of the quadratic systems free of movable critical points (Theorem 1). Here, $n\in\mathbf{Z}$, $a$, $b$, etc. are functions of the independent variable, $a_{0}$ and $b_{0}$ are constants.
The labeling of the systems in Table 1 parallels Bureau’s in [Bur81b] and [Bur81c]. Systems $\text{V}_{0}$, $\text{VII}_{0}$, $\text{IX.A}_{0}(n)$ and $\text{IX.B}_{0}(n)$ in Table 1 are absent from Bureau’s classification. They are simple equations, which are very easily integrated, but their absence challenges Bureau’s claim of comprehensiveness in the presentation of Table III in [Bur81c, §28]: “Il nous paraît utile de réunir dans le tableau suivant les représentants de tous les s.d. [systèmes différentiels] polynomiaux stables de degré deux” (we thought it useful to gather representatives of all stable polynomial differential systems of degree two in the following table). We give slightly simpler formulas than Bureau’s for systems I, III and IV, and correct an error in the analysis of system IX.A(3). Our normal forms for systems XII, XIII and XIV differ from Bureau’s; they permit to relate them in a simpler way to relevant second-order equations in one variable, and to express in a more concise manner the conditions for the absence of movable critical points. We take into account the symmetries of systems XII and XIII, which allows us to avoid the repetition of some calculations (or, through it, to double or triple check their correctness). Apart from this, the results summarized in Table 1 are due to Bureau.
Some systems in Table 1 are birationally equivalent to the first-order system in two variables associated a Painlevé equation, and, through these identifications, some pairs of systems belonging to different rows in Table 1 can be shown to be birationally equivalent. We have the following result, obtained in the course of establishing Theorem 1.
@@
For the systems in Table 1 that are solved by Painlevé transcendents, a bimeromorphic map $\phi:\mathbf{C}^{2}\times\mathbf{C}\dashrightarrow\mathbf{C}^{2}\times\mathbf{C}$, of the form $(\zeta,t)\mapsto(\xi(\zeta,t),t)$, with $\xi$ birational as a function of $\zeta$ for most values of $t$, transforms them into the first-order system in two variables associated to the corresponding Painlevé equation.
Painlevé I’s birational class:
The Painlevé I equation and some of its degenerations are given by the second-order equation
$u^{\prime\prime}=6u^{2}+f,\;\text{with }f^{\prime\prime}=0.$ (2)
Systems V, IX.B(2), IX.B(5) and XIV are birationally equivalent to the first-order system in two variables associated to it (and, in particular, are birationally equivalent to one another). More precisely, we have the following.
V:
System V is exactly the first-order system in two variables associated to (2), for the same parameter $f$.
IX.B(2):
From a solution $(y,z)$ to the system IX.B(2) with parameter $q$,
$u={\textstyle\frac{1}{6}}z+q$
is a solution to equation (2) with parameter $f=q^{\prime\prime}-6q^{2}$. (Since $u$ is polynomially expressed in terms of $y$ and $z$, and the system solved by $y$ and $z$ is polynomial in $y$ and $z$, we have that $u^{\prime}$ may be polynomially expressed in terms of $y$ and $z$). Reciprocally, for the equation (2) with parameter $f$, and a solution $u$ to it, for $q$ defined by $q^{\prime\prime}=6q^{2}+f$,
$(y,z)=\left(\frac{u^{\prime}-q^{\prime}}{u-q},6(u-q)\right)$
is a solution to system IX.B(2) with parameter $q$. These two transformations are inverses of one another.
IX.B(5):
From a solution $(y,z)$ to the system IX.B(5) with parameter $q$,
$u=y^{2}-{\textstyle\frac{1}{3}}z-2q,$
is a solution to equation (2) with parameter $f=q^{\prime\prime}-6q^{2}$. Reciprocally, from a solution $u$ to equation (2) with parameter $f$, for $q$ defined by $f=q^{\prime\prime}-6q^{2}$,
$(y,z)=\left(-\frac{1}{2}\left(\frac{u^{\prime}-q^{\prime}}{u-q}\right),\frac{3}{4}\left(\frac{u^{\prime}-q^{\prime}}{u-q}\right)^{2}-3u-6q\right)$
is a solution to system IX.B(5) with this parameter $q$. These two transformations are inverses of one another.
XIV:
@@
Consider a solution $(y,z)$ to system XIV with parameters $(r,p)$, and let $q=\frac{1}{12}(p^{\prime}+p^{2}-r)$. Then,
$u={\textstyle\frac{1}{6}}zy+{\textstyle\frac{1}{6}}r+q$
is a solution to equation (2) with parameter $f=q^{\prime\prime}-6q^{2}$. Reciprocally, from a solution $u$ to equation (2) with parameter $f$, for $q$ defined by $q^{\prime\prime}-6q^{2}=f$, and $p$ and $r$ satisfying the conditions of system XIV with respect to $q$,
$(y,z)=\left(\frac{6u-6q-r}{\frac{u^{\prime}-q^{\prime}}{u-q}-p},\frac{u^{\prime}-q^{\prime}}{u-q}-p\right)$
is a solution to system XIV with parameters $(r,p)$. These two transformations are inverses of one another.
Painlevé II’s birational class:
The Painlevé II equation and some of its degenerations are given by the second-order equation
$u^{\prime\prime}=2u^{3}+fu+\alpha,\;\text{with }\alpha\in\mathbf{C},f^{\prime\prime}=0.$ (3)
@@
Systems IX.B(3) and XIII are birationally equivalent to the first order system in two variables associated to it.
IX.B(3):
From a solution $(y,z)$ to system IX.B(3) with parameters $(f,a_{0})$, $y$ is a solution to (3) for the same $f$ and $\alpha=a_{0}-\frac{1}{2}f^{\prime}$. Reciprocally, from a solution $u$ to the equation (3) with parameters $(f,\alpha)$,
$(y,z)=\left(u,u^{\prime}+u^{2}+{\textstyle\frac{1}{2}}f\right)$
is a solution to the system IX.B(3) whose parameters are the same $f$ and $a_{0}=\alpha+\frac{1}{2}f^{\prime}$. These two nonlinear polynomial transformations are inverses of one another.
XIII:
For a solution $(y,z)$ of system XIII with parameters $(f,p)$,
$u={\textstyle\frac{1}{2}}y-p$
is a solution to equation (3) for the same $f$ and
$\alpha=-(p^{\prime\prime}-2p^{3}-fp).$ (4)
Reciprocally, given $f$ with $f^{\prime\prime}=0$ and $\alpha\in\mathbf{C}$, from a solution $u$ to the associated system (3), for $p$ defined by (4),
$(y,z)=\left(2(u+p),u-p+\frac{u^{\prime}+p^{\prime}}{u+p}\right)$
is a solution to system XIII with parameters $(f,p)$. These two transformations are inverses of one another.
Painlevé IV’s birational class:
The Painlevé IV equation and some of its degenerations are given, in Gambier’s normalization [Gam10], by
$u^{\prime\prime}=\frac{(u^{\prime})^{2}}{2u}+\frac{3}{2}u^{3}+4fu^{2}+2(f^{2}-\alpha)u+\frac{\beta}{u},\;\text{with }\alpha,\beta\in\mathbf{C},f^{\prime\prime}=0.$ (5)
System XII is birationally equivalent to the first-order system in two variables coming from it. If $(y,z)$ is a solution to system XII, $z$ solves Eq. (5) with the same $f$ and
$\alpha={\textstyle\frac{1}{2}}a_{0}+b_{0}-f^{\prime},\;\beta=-{\textstyle\frac{1}{2}}a_{0}^{2}.$ (6)
Reciprocally, from a solution $u$ to equation (5) with parameters $(f,\alpha,\beta)$, for constants $a_{0}$ and $b_{0}$ satisfying (6),
$(y,z)=\left(-\frac{u^{\prime}}{2u}+\frac{u}{2}+f-\frac{a_{0}}{2u},u\right)$
is a solution to system XII with these parameters.
@@
Systems IX.B(3) and XIV precede Bureau’s investigations, and already appear in Malmquist’s formulation of Painlevé’s equations II and IV as Hamiltonian systems [Mal23, §33]. The first-order systems directly associated to these equations are not Hamiltonian, but the birational models given by the quadratic systems IX.B(3) and XIV are. This is presented in Table 2, along with the Hamiltonian of the first Painlevé equation, directly related to system V.
Equation Hamiltonian system
Painlevé I $\frac{1}{2}z^{2}-2y^{3}-fy$ V
Painlevé II $y^{2}z+\frac{1}{2}z^{2}-\frac{1}{2}fz-a_{0}y$ IX.B(3)
Painlevé IV $y^{2}z-yz^{2}-2fyz+a_{0}y+b_{0}z$ XII
Table 2. Malmquist’s Hamiltonian formulation for the Painlevé I, II and IV equations. Here, $a_{0},b_{0}\in\mathbf{C}$, $f^{\prime\prime}=0$.
The less satisfactory part of the classification concerns system VIII, the only one for which the necessary and sufficient conditions for the absence of movable critical points do not explicitly appear, neither in Bureau’s work, nor in Table 1. These systems form an infinite family, indexed by a positive integer $n$, and the complexity of a certain necessary condition for the absence of movable critical points grows with $n$; it seems difficult to give an explicit expression for this condition for every $n$. At the same time, the equations are, conceptually, very simple, as they factor through Riccati equations. In any case, there remains the question of whether this necessary condition is also a sufficient one. We will show that this is indeed the case, by studying the birational geometry of these systems, in the spirit of Okamoto’s work on the spaces of initial conditions of the Painlevé equations [Oka79], which is, on its turn, motivated by Painlevé’s 17th of his Leçons de Stockholm [Pai97]. This will be done in Section 3.6.
@@
No system X appears in Table 1. A system X does appear in Bureau’s Table III in [Bur81c], but there are several issues with it. First, in Bureau’s rather structured text, the discussion of any system labeled X would be expected to appear in [Bur81b] after §20 but before §21, and is nowhere to be found. Second, the system appearing in Bureau’s table is $y^{\prime}=yz+a$, $z^{\prime}=(n-1)yz+b$, with $n>1$; however, the parameter $n$ in it is rendered irrelevant by scaling $y$, which suggests that there is a misprint in the formula. Indeed, as per the caption for this system in Bureau’s table, the quadratic part of the system should be the one in formula (19.3) in [Bur81a], which is said to be equivalent to the quadratic system (19.2) there, thus belonging to system IX. We suspect that our systems $\text{IX.A}_{0}(n)$ and $\text{IX.B}_{0}(n)$ play the role conceived by Bureau for his system X. In any case, the classification in Theorem 1 has no need for this system.
@@ 2.1. Quadratic homogeneous equations and univalence
The limit as $\alpha\to 0$ is the autonomous, homogeneous equation $\mathrm{d}\xi/\mathrm{d}\tau=P_{n}(\xi,t_{0})$. If the original equation is free of movable critical points, the latter will have only single-valued solutions [Pai00] (see also [Bur80]). Thus, in our setting, the first problem is to determine the equations of the form
@@ V $\begin{array}[]{l}y^{\prime}=0\\ z^{\prime}=6y^{2}\end{array}$
VII $\begin{array}[]{l}y^{\prime}=0\\ z^{\prime}=yz\end{array}$ $\infty$
@@
$\begin{array}[]{l}y^{\prime}=y[y-(n+2)z]\\ z^{\prime}=-z(ny+z)\end{array}$ 1,n+1,n11,n+1,-n-1
@@ XI $\begin{array}[]{l}y^{\prime}=y(y-z)\\ z^{\prime}=z(z-y)\end{array}$ $2,2,\infty$
XII $\begin{array}[]{l}y^{\prime}=y(y-2z)\\ z^{\prime}=z(z-2y)\end{array}$ $3,3,3$
@@ XIV $\begin{array}[]{l}y^{\prime}=y(2z-y)\\ z^{\prime}=z(y-z)\end{array}$ $2,3,6$
Table 3. Representatives of the univalent quadratic homogeneous systems, with the indices of their simple radial orbits (we adhere to Bureau’s normalization except in system XII, where our normal form differs from Bureau’s by a change of sign in $z$).
@@
$\frac{s^{\prime\prime}}{(s^{\prime})^{2}}=F(s),$ (8)
for some rational function $F$ (a similar equation is satisfied by $1/s$). Briot and Bouquet gave necessary and sufficient conditions on $F$ for the univalence of a function $s$ satisfying such an equation [BB55]: at the finite values of $s$, $F$ has only simple poles, and their residues are either equal to 1-1, or are of the form 1/k11/k-1 for $k\in\mathbf{Z}$, plus the analogous condition at $s=\infty$ coming from the equation satisfied by $1/s$ (see also Painlevé [Pai00, §10]; see [Gui06, Section 2.2] for an interpretation in terms of affine structures). The index of the pole of (8) is the associated integer $k$ (we will say that a simple pole with residue 1-1 has index $\infty$). We refer the reader to the closing remarks in [Con99, §2.3] for a discussion around this terminology.
A radial orbit for a homogeneous equation of the form (7) is a line through $(0,0)$ that is invariant by the equation (that is a union of solutions). If not every line through the origin is a radial orbit, the equation has three of them (counted with multiplicity). The poles of the associated equation (8) come from radial orbits, although some radial orbits may fail to produce them. The index of a radial orbit is the one of the corresponding pole in equation (8); a simple radial orbit which produces no poles for (8) is said to have index $1$. If an equation has three radial orbits, with indices $n_{i}$,
@@
The classification of the quadratic homogeneous equations (7) with three different radial orbits that have only univalent solutions amounts to solving this equation as a Diophantine one (admitting that $n_{i}$ can take the value $\infty$). The last six lines of Table 3 are obtained in this way (the linear equivalence class of such an equation is determined by its indices, and the equations thus obtained happen to be univalent). The other lines of this table follow from organizing quadratic homogeneous vector fields according to the multiplicities of their radial orbits, to the indices of the simple ones, and from imposing the conditions guaranteeing the simplicity of the poles of the associated equation (8).